# Complexity in Complexity: Understanding Visual Complexity Through Structure, Color, and Surprise

### v2

@@ Abstract
Understanding human perception of visual complexity is crucial in visual cognition. Recently (Shen, et al. 2024) proposed an interpretable segmentation-based model that accurately predicted complexity across various datasets, supporting the idea that complexity can be explained simply. In this work, we investigate the failure of their model to capture structural, color and surprisal contributions to complexity. To this end, we propose Multi-Scale Sobel Gradient which measures spatial intensity variations, Multi-Scale Unique Color which quantifies colorfulness across multiple scales, and surprise scores generated using a Large Language Model. We test our features on existing benchmarks and a novel dataset containing surprising images from Visual Genome. Our experiments demonstrate that modeling complexity accurately is not as simple as previously thought, requiring additional perceptual and semantic factors to address dataset biases. Thus our results offer deeper insights into how humans assess visual complexity.
@@
11footnotetext: These authors contributed equally to the supervision of the project.
@@ Methods
In this work, we evaluate the effectiveness of sets of features to explain visual complexity on datasets using linear regression models. We follow the procedure in (?, ?), and fit $M$ repetitions of 3-fold cross-validated linear regression. $M$ is determined by the size of the dataset to address statistical variability in smaller datasets. We measure performance using the mean Spearman correlation coefficient across all test splits. We compare the performance of our models to the same baselines in (?, ?): handcrafted features (?, ?, ?) and deep learning approaches (?, ?, ?) for comparison. The code for our analysis, experimental setup, and dataset will be made publicly available upon acceptance.
@@
To the best of our knowledge, there is currently no existing image set with surprising images that can be used to study visual complexity. To fill this gap, we introduce a new dataset containing 100 highly surprising images and 100 (on average, less surprising) images, the latter of which were sampled uniformly at random from the Visual Genome, subject to stratification according to bins of complexity based on IC9600 scores (?, ?). To confirm that our subjectively selected images are more surprising than randomly sampled ones, we compare the distributions of LLM-generated surprise scores between the two subsets. A histogram of these scores shows clear separation: sampled images cluster at lower values, while handpicked images have higher surprise scores (mean: $39.45$ for ordinary, $72.35$ for surprising). A Kolmogorov-Smirnov test confirms a significant difference ($D=0.67,p<0.001$), supporting our hypothesis.
We collect visual complexity ratings from humans using an online experiment on Prolific (?, ?). Participants are shown pairs of images and asked to select the one that is more visually complex, similarly to the procedure in (?, ?). We followed the sampling strategy in (?, ?), and collect $6000=200\times 30$ comparisons by sampling images with probabilities inversely proportional to the total number of times they have already been selected. We ensure that each pairwise comparison is assessed by three different participants. Hence there are in total $18,000$ evaluations. Each participant evaluates $200$ pairwise comparisons in one session. We convert pairwise comparisons into scalar complexity ratings using the Bradley-Terry algorithm (?, ?).
@@
We propose the Multi-Scale Sobel Gradient (MSG) algorithm (Algorithm 1) to capture structure in images. MSG analyzes spatial intensity variations across multiple resolutions in an RGB image. Note that the Sobel operator is typically applied to grayscale images (?, ?), but we found in ablations that the grayscale version of the algorithm (which first converts the colored images to grayscale) did not perform as well. Hence we use the color version of the algorithm for the rest of this work.
@@
Generating Surprise Scores using LLMs
Algorithm 3 Zero-shot-CoT
Provide your reasoning and numeric rating as follows:
Reasoning: [your explanation]
[…]
@@ Results
To explore quantitatively how well our gemini-1.5-flash surprise scores aligned with human judgments of surprise, we computed Spearman’s correlation between gemini-1.5-flash scores and the human surprise scores we collected. We find a high correlation of 0.75. This supports the fact gemini-1.5-flash explanations are good proxies for human judgments of image surprisal which contribute to human perceptions of visual complexity. In summary, gemini-1.5-flash surprise scores improve the predictive power of our complexity model on the SVG dataset and the natural language explanations generated alongside these scores facilitate a clearer understanding of why the image is complex.
The last row of Table 1 shows the performance of our penultimate model including MSG, MUC, surprise as well as the two features from (?, ?). Interestingly, we find combining MSG and MUC with surprise only improves performance on all datasets over using either of these features alone. Furthermore, our penultimate model is significantly more performant than every baseline except (?, ?)’s supervised network which is competitive. Taken together, in contrast to what (?, ?) previously suggested, the complexity of naturalistic images cannot be explained simply using two generic segment and object features. Some datasets require additional features to explain visual complexity fully.
Refer to caption (a) $\sqrt{\textit{num\_seg}}$ = 191 $\sqrt{\textit{num\_class}}$ = 9 MSG = 59 MUC = 922 Surprise = 25 Refer to caption (b) $\sqrt{\textit{num\_seg}}$ = 172 $\sqrt{\textit{num\_class}}$ = 9 MSG = 56 MUC = 820 Surprise = 85 Surprise as the hidden dimension We present a case where both images share similar visual features, including segmentations, classes, MSG, and MUC. The right image, with lower feature values, was initially predicted to have lower complexity. The left image received a complexity score of 60, while the right received 55. However, the element of surprise shifts the outcome, with improved predictions at 52 (72) compared to ground truth values of 25 (74). Despite their similarity, humans may perceive the second image as more complex due to its unusual composition, where multiple traffic signs are clustered together.
@@ Discussion
Patch symmetry (?, ?, ?, ?) and edge density (?, ?, ?) are two commonly used low-level features for capturing structural information in an image. We run ablation studies to investigate whether MSG is more performant than these two features. We evaluate on all 16 datasets: VISC, RSIVL, SAVOIAS (5 categories - excluding Advertisement and Visualizations due to heavy presence of text), IC9600 (8 categories), and SVG using permutation tests. We compare MSG against patch symmetry because (?, ?) reported that it captures structural information well on the VISC dataset. We compare MSG against Canny edge density because Canny is a standard metric for predicting visual complexity. It has been applied in (?, ?, ?, ?).
MSG performs significantly better than Canny edge density in five datasets. For example, on IC9. Abstract, MSG achieves significantly higher correlation with the complexity compared to edge density (0.517 vs 0.705). Edge density only outperformed MSG on two datasets. Compared to patch symmetry, MSG performs better on 7 datasets. Patch symmetry outperforms MSG on only one dataset. Our ablations show that MSG is currently the most performant structural feature for predicting complexity, outperforming both edge density and patch symmetry across diverse image datasets.
@@
We perform ablation studies to investigate the necessity of multi-scale computations for MSG and MUC. We measure the correlation between complexity and our visual features using both single and multiple scales. We find that multi-scale MSG outperforms single-scale MSG on 6 datasets, while single-scale MSG outperforms multi-scale MSG on just two datasets out of the 16 datasets tested. The advantage of the multi-scale approach was particularly pronounced in Sav. Suprematism, where correlation increased substantially from 0.474 to 0.616. Similarly, multi-scale MUC outperformed single-scale MUC on 5 datasets, while single-scale MUC outperformed multi-scale MUC on only one dataset. The multi-scale versions of MSG and MUC maximize the correlation with the perceived visual complexity.
@@
The existing literature on visual complexity has largely concentrated on identifying low-level image features that define complexity, rather than examining its perceptual implications for the observer. However, as noted by (?, ?), complexity inherently relates to the difficulty of cognitive processing, suggesting that surprising images should be perceived as more complex. We provide first evidence in support of this hypothesis, showing that surprise, as determined by a LLM, contributes significantly to visual complexity on our SVG dataset.
Prior work links image memorability, a closely related concept to complexity (?, ?), to distinctiveness, finding that unique or atypical features enhance retention. (?, ?, ?) find that distinct facial features are more memorable. (?, ?) find that images that deviate from their neighbors in a CNN embedding space have increased recall. We observe a weak positive correlation between human surprisal ratings and both AMNet-predicted ($\rho=0.25,p<0.05$) (?, ?) and ResMem-predicted memorability ($\rho=0.25,p<0.05$) (?, ?), suggesting that surprise may contribute to, or even be a common cause for both judgments of complexity and memorability. Future works should elucidate the relationship between complexity, memorability and surprise.

### v3

@@ Abstract
Understanding how humans perceive visual complexity is a key area of study in visual cognition. Previous approaches to modeling visual complexity assessments have often resulted in intricate, difficult-to-interpret algorithms that employ numerous features or sophisticated deep learning architectures. While these complex models achieve high performance on specific datasets, they often sacrifice interpretability, making it challenging to understand the factors driving human perception of complexity. Recently (Shen, et al. 2024) proposed an interpretable segmentation-based model that accurately predicted complexity across various datasets, supporting the idea that complexity can be explained simply. In this work, we investigate the failure of their model to capture structural, color and surprisal contributions to complexity. To this end, we propose Multi-Scale Sobel Gradient (MSG) which measures spatial intensity variations, Multi-Scale Unique Color (MUC) which quantifies colorfulness across multiple scales, and surprise scores generated using a Large Language Model. We test our features on existing benchmarks and a novel dataset (Surprising Visual Genome) containing surprising images from Visual Genome. Our experiments demonstrate that modeling complexity accurately is not as simple as previously thought, requiring additional perceptual and semantic factors to address dataset biases. Our model improves predictive performance while maintaining interpretability, offering deeper insights into how visual complexity is perceived and assessed. Our code, analysis and data are available at our Github repository [^1] https://github.com/Complexity-Project/Complexity-in-Complexity.
@@
$\ast$\$$\ast$\$footnotetext: Joint supervision
@@ Methods
In this work, we evaluate the effectiveness of sets of features to explain visual complexity on datasets using linear regression models. We follow the procedure in (?, ?), and fit $M$ repetitions of 3-fold cross-validated linear regression. $M$ is determined by the size of the dataset to address statistical variability in smaller datasets. We measure performance using the mean Spearman correlation coefficient across all test splits. We compare the performance of our models to the same baselines in (?, ?): handcrafted features (?, ?, ?) and deep learning approaches (?, ?, ?).
@@
To the best of our knowledge, there is currently no existing image set with surprising images that can be used to study visual complexity. To fill this gap, we introduce a new dataset containing 100 highly surprising images and 100 (on average, less surprising) images, the latter of which were sampled uniformly at random from the Visual Genome, subject to stratification according to bins of complexity based on IC9600 scores (?, ?). To confirm that our subjectively selected images are more surprising than randomly sampled ones, we compare the distributions of LLM-generated surprise scores (as explained in Section Generating Surprise Scores using LLMs) between the two subsets. A histogram of these scores shows clear separation: sampled images cluster at lower values, while handpicked images have higher surprise scores (mean: $39.45$ for ordinary, $72.35$ for surprising). A Kolmogorov-Smirnov test confirms a significant difference ($D=0.67,p<0.001$), supporting our hypothesis.
We collect visual complexity ratings from humans using an online experiment on Prolific (?, ?). Participants are shown pairs of images and asked to select the one that is more visually complex, similarly to the procedure in (?, ?). We followed the sampling strategy in (?, ?), and collect $6000=200\times 30$ comparisons by sampling images with probabilities inversely proportional to the total number of times they have already been selected. We ensure that each pairwise comparison is assessed by three different participants. Hence there are in total $18,000$ evaluations. Each participant evaluates $200$ pairwise comparisons in one session. We also inserted attention checks into the experiment to ensure participant engagement and data reliability. We convert pairwise comparisons into scalar complexity ratings using the Bradley-Terry algorithm (?, ?). Initially, the final ratings were mostly clustered around zero, a phenomenon also observed in (?, ?). To address this, we applied the same procedure as in their work. Before using the Bradley-Terry model, we rescaled the initial probability matrix from the range $[0,1]$ to $[0.33,0.66]$. This adjustment helped distribute the final ratings more evenly within a narrower and more interpretable range. We also collect human surprise ratings to validate our LLM-generated scores using a similar experiment.
Generating Surprise Scores using LLMs
Algorithm 1 Zero-shot-CoT
Provide your reasoning and numeric rating as follows:
Reasoning: [your explanation]
[…]
@@
We propose the Multi-Scale Sobel Gradient (MSG) algorithm (Algorithm 2) to capture structure in images. MSG applies the Sobel operator across multiple resolutions to RGB images, functioning as an asymmetry detector within $k\times k$ patches, where $k$ represents the kernel size. For symmetric patches, the left and right columns (or rows in horizontal application) of the kernel counterbalance each other, resulting in lower MSG values. Consequently, images exhibiting greater patch-level symmetry produce smaller MSG measurements. Note that the Sobel operator is typically applied to grayscale images (?, ?), but we found in ablations that the grayscale version of the algorithm (which first converts the colored images to grayscale) did not perform as well. Based on our permutation tests across 16 datasets, the colored version demonstrated statistically significant superiority over the grayscale version in 6 datasets, while the grayscale version outperformed the colored in 2 datasets. The remaining 8 datasets showed no statistically significant differences between the two versions. Hence we use the color version of the algorithm for the rest of this work.
@@ Results
Two examples of explanations generated from gemini-1.5-flash using CoT prompting (?, ?, ?), can be seen in Figure 2. In the left image, gemini-1.5-flash successfully identifies the intuitively surprising fact that sheep are racing, an unusual activity for sheep. In the right image, gemini-1.5-flash identifies that it’s unusual for the man to be doing an uncontrolled backflip, arguably consistent with human intuition for why the image is surprising.
@@
To explore quantitatively how well our gemini-1.5-flash surprise scores aligned with human judgments of surprise, we computed Spearman’s correlation between gemini-1.5-flash scores and the human surprise scores we collected. The results revealed a strong positive correlation of 0.73, indicating substantial agreement between model-generated and human-derived assessments of visual surprise. This supports the fact gemini-1.5-flash explanations are good proxies for human judgments of image surprisal which contribute to human perceptions of visual complexity. In summary, gemini-1.5-flash surprise scores improve the predictive power of our complexity model on the SVG dataset and the natural language explanations generated alongside these scores facilitate a clearer understanding of why the image is complex.
The last row of Table 1 shows the performance of our final model including MSG, MUC, surprise as well as the two features from (?, ?). Interestingly, we find combining MSG and MUC with surprise only improves performance on all datasets over using either of these features alone. Furthermore, our final model is significantly more performant than every baseline except (?, ?)’s supervised network which is competitive. Taken together, in contrast to what (?, ?) previously suggested, the complexity of naturalistic images cannot be explained simply using two generic segment and object features. Some datasets require additional features to explain visual complexity fully.
Lastly, we evaluate the performance of gemini-1.5-flash on our SVG dataset using two distinct prompting strategies. In the first approach, the model estimated visual complexity without an explicit definition, replicating the open-ended methodology from our Prolific human-subject experiments. In the second approach, we provided the canonical definition of visual complexity as “the level of detail or intricacy contained within an image” (?, ?) to anchor the model’s judgments. Both strategies leveraged Chain-of-Thought reasoning to generate structured rationales from gemini-1.5-flash. The resulting complexity ratings showed strong internal consistency (Spearman’s $\rho=0.89$ between the two strategies) and high alignment with human ground-truth ratings ($\rho=0.83$ for both). Notably, our final integrated model—which combines pixel-level features, object-level segmentation metrics, and surprise scores from gemini-1.5-flash (quantifying how surprising an image is)—outperformed gemini-1.5-flash ’s standalone visual complexity evaluations on the SVG dataset.
Refer to caption (a) num_seg = 191 num_class = 9 MSG = 59 MUC = 922 Surprise = 25 Refer to caption (b) num_seg = 172 num_class = 9 MSG = 56 MUC = 820 Surprise = 85 Surprise as the hidden dimension We present a case where both images share similar visual features, including segmentations, classes, MSG, and MUC. The right image, with lower feature values, was initially predicted to have lower complexity. The left image received a complexity score of 60, while the right received 55. However, the element of surprise shifts the outcome, with improved predictions at 52 (72) compared to ground truth values of 25 (74). Despite their similarity, humans may perceive the second image as more complex due to its unusual composition, where multiple traffic signs are clustered together.
@@ Discussion
Numerosity Perception
As the baseline approach, (?, ?) uses the square root of the number of segmentations and classes, an empirical finding that demonstrated better performance than using the raw values. One reason why the square root works better than the raw values can be explained by the Weber-Fechner Law (?, ?), which establishes that subjective perception typically relates to stimulus intensity in a nonlinear manner. According to this principle, the just-noticeable difference between two stimulus intensities is proportional to the absolute stimulus intensity. From a practical perspective, this means participants can more readily distinguish between images containing 10 versus 20 segmentations than between those with 110 versus 120 segmentations, despite the absolute difference being identical. We also tested applying logarithmic transformation to the number of segmentations and classes, but the square-root transformation achieved slightly better results in the regression model, so we decided to stick with it.
@@
Patch symmetry (?, ?, ?, ?) and edge density (?, ?, ?) are two commonly used low-level features for capturing structural information in an image. We run ablation studies to investigate whether MSG is more performant than these two features. We evaluate on all 16 datasets: VISC, RSIVL, SAVOIAS (5 categories - excluding Advertisement and Visualizations due to heavy presence of text), IC9600 (8 categories), and SVG using permutation tests. We compare MSG against patch symmetry because (?, ?) reported that it captures structural information well on the VISC dataset. We compare MSG against Canny edge density because Canny is a standard metric for predicting visual complexity. It has been applied in (?, ?, ?, ?). MSG performs significantly better than Canny edge density in five dataset (IC9. Abstract, IC9. Paintings, Sav. Scenes, IC9. Advertisement, and Sav. Interior Design). For example, on IC9. Abstract, MSG achieves significantly higher correlation with the complexity compared to edge density (0.517 vs 0.705). Edge density only outperformed MSG on two datasets (IC9. Person and Sav. Art). Compared to patch symmetry, MSG performs better on 7 datasets (IC9. Abstract, IC9. Paintings, IC9. Scenes, IC9. Advertisement, IC9. Architecture, IC9. Person, and Sav. Art). Patch symmetry outperforms MSG on only one dataset (Sav. Interior Design dataset). Our ablations show that MSG is currently the most performant structural feature for predicting complexity, outperforming both edge density and patch symmetry across diverse image datasets.
@@
We perform ablation studies to investigate the necessity of multi-scale computations for MSG and MUC. We measure the correlation between complexity and our visual features using both single and multiple scales. By single scale, we refer to the application of our algorithms on the whole image only once ($W={1.0}$ and $S={1}$) without rescaling. We find that multi-scale MSG outperforms single-scale MSG on 8 datasets (VISC, IC9 Abstract, Sav. Objects, Sav. Scenes, Sav. Suprematism, IC9 Transport, IC9 Objects, and SVG), while single-scale MSG outperforms multi-scale MSG on just two datasets (IC9 Paintings and IC9 Architecture) out of the 16 datasets tested. The advantage of the multi-scale approach was particularly pronounced in Sav. Suprematism, where correlation increased substantially from 0.474 to 0.616. Similarly, multi-scale MUC outperformed single-scale MUC on 5 datasets (IC9. Abstract, IC9. Paintings, Sav. Art, IC9. Advertisement, IC9. Architecture), while single-scale MUC outperformed multi-scale MUC on only one dataset (IC9 Transport). The multi-scale versions of MSG and MUC maximize the correlation with the perceived visual complexity.
@@
The existing literature on visual complexity has largely concentrated on identifying low-level image features that define complexity, rather than examining its perceptual implications for the observer. However, as noted by (?, ?), complexity inherently relates to the difficulty of cognitive processing, suggesting that surprising images should be perceived as more complex. We provide evidence in support of this hypothesis, showing that surprise, as determined by an LLM, contributes significantly to visual complexity on our SVG dataset.
Prior work links image memorability, a closely related concept to complexity (?, ?), to distinctiveness, finding that unique or atypical features enhance retention. (?, ?, ?) find that distinct facial features are more memorable. (?, ?) find that images that deviate from their neighbors in a CNN embedding space have increased recall. We observe a weak positive correlation between human surprisal ratings and both AMNet-predicted ($\rho=0.25,p<0.05$) (?, ?) and ResMem-predicted memorability ($\rho=0.25,p<0.05$) (?, ?), suggesting that surprise may contribute to, or even be a common cause for both judgments of complexity and memorability. Future works should elucidate the relationship between complexity, memorability, and surprise.
@@
Appendix
Results on All Datasets
Table 2 compares the performance of the baseline approach, our final model, and Feng’s supervised model across all datasets.
Table 2: Correlations with the ground truth perceived complexity ratings. Bold indicates the best model performance for each dataset. Improvements from the baseline are indicated with arrows.
Dataset Baseline Our final model Feng (sup.)
VISC 0.56 0.71 $\uparrow 0.15$ 0.72
RSIVL 0.83 0.84 $\uparrow 0.01$ 0.83
SVG 0.78 0.85 $\uparrow 0.07$ 0.83
Sav. Art 0.73 0.82 $\uparrow 0.09$ 0.82
Sav. Sup. 0.89 0.94 $\uparrow 0.05$ 0.84
Sav. Int. 0.61 0.87 $\uparrow 0.26$ 0.89
Sav. Objects 0.82 0.85 $\uparrow 0.03$ 0.83
Sav. Scenes 0.77 0.81 $\uparrow 0.04$ 0.87
IC9. Arch. 0.66 0.77 $\uparrow 0.11$ 0.93
IC9. Abstract 0.66 0.84 $\uparrow 0.18$ 0.95
IC9. Scenes 0.77 0.83 $\uparrow 0.06$ 0.93
IC9. Objects 0.80 0.84 $\uparrow 0.04$ 0.92
IC9. Transport 0.74 0.79 $\uparrow 0.05$ 0.88
IC9. Paintings 0.83 0.87 $\uparrow 0.04$ 0.95
IC9. Adv. 0.75 0.82 $\uparrow 0.07$ 0.93
IC9. Person 0.59 0.66 $\uparrow 0.07$ 0.85
[…]
When evaluating on datasets not included in Feng’s training data, our model outperforms Feng’s supervised approach on 4 out of 8 datasets, while Feng’s model performs better on 3 datasets (with one dataset resulting in a tie), giving us a final record of 4/3/1.
Feng’s supervised method shows superior performance across all IC9 datasets as expected since their model was specifically trained on this data distribution (in-domain evaluation). In contrast, our approach maintains strong performance while using interpretable features that don’t require training on any dataset.
Additional Experiment Details
In our main paper, we briefly described the creation of our Surprising Visual Genome (SVG) dataset and our methodology for collecting human complexity ratings. Here, we provide additional details about our experimental procedure.
The visual complexity assessment task was conducted using Prolific (?, ?) as the participant recruitment platform. Each participant completed 200 pairwise comparisons in a single session. The study, which took approximately 15 minutes to complete, offered payment of 2.61 pounds. All participants were recruited from the UK.
The experiment was implemented using jsPsych (?, ?), a JavaScript library for creating behavioral experiments in web browsers. After reading instructions and completing practice trials, participants were presented with image pairs and asked to select the one they perceived as more visually complex (see Figure 8). Initial instructions are provided in Frame 7.
Thank you for your consent to participate in our study. In this study, you will be shown pairs of images and asked to judge which one appears more visually complex to you.
You will make around 200 comparisons in total. For each pair, simply click on the image you judge to be more surprising. There are no right or wrong answers - we are interested in your perception of surprise.
At the end of the study, we will ask brief questions to better understand your judgements of visual complexity. The complete study will take approximately 15 minutes.
We will now show you some example pairs to familiarize you with the task. Click “Start” to begin.
Figure 7: Initial instructions provided to participants at the beginning of the visual complexity experiment
Figure 8: Interface of the pairwise comparison experiment for visual complexity assessment. Participants were instructed to select the image they perceived as more visually complex by clicking on it. The number of comparisons remaining was displayed in the bottom right corner.
To ensure data quality, we embedded three attention check trials among the 200 comparisons. In these trials, participants were shown identical images, except one image contained a text overlay with explicit instructions to select it. Participants who failed more than one attention check were excluded from the analysis and replaced with new participants to maintain our target sample size.
After completing all comparisons, participants responded to questions about their strategy for rating visual complexity, whether they found particular types of images consistently more complex, and if they had additional comments about their experience. These responses provided qualitative insights into the factors influencing complexity judgments, as highlighted in the main paper.
The pairwise comparison data was processed using the Bradley-Terry algorithm (?, ?) to convert the binary choices into continuous complexity scores. To address the issue of scores clustering around zero, we applied the rescaling procedure described in (?, ?), which involved transforming the probability matrix from [0,1] to [0.33, 0.66] before applying the Bradley-Terry model.
To collect human surprise ratings, we utilized the same experimental setting with minor modifications to the instructions and text to focus on surprisal rather than complexity.
Our complete experimental materials, including stimuli, code for the web-based interface, data processing scripts, and analysis pipelines, are available at https://github.com/Complexity-Project/Experiment to facilitate replication and extension of our work.
SVG dataset
To curate the SVG dataset, we employed a targeted approach, first manually identifying candidates with high surprisal value from the Visual Genome dataset, then validating our selections using an automated method (using an LLM). We considered an alternative strategy—directly evaluating surprise values across the entire corpus using an LLM and selecting those exceeding a threshold—but determined this was impractical for the following: Identifying truly distinctive surprising images, as exemplified in Figure 9 and subsequently confirmed by participant responses in Figure 4, required examining a substantial portion of the dataset, which consists of $\geq 100$K images. The computational demands of LLM processing at this scale imposed substantial practical constraints: either (1) lengthy processing times due to API rate limits, or (2) prohibitive costs for premium API access that would bypass these limitations. Therefore, we implemented a hybrid approach, first manually selecting candidate images with high surprisal potential, followed by LLM verification to ensure selection validity.
(a) G: 15, H: 4, C: 5
(b) G: 35, H: 32, C: 81
(c) G: 65, H: 66, C: 27
(d) G: 75, H: 76, C: 56
(e) G: 85, H: 98, C: 72
(f) G: 85, H: 85, C: 81
Figure 9: Representative samples from the SVG dataset. G indicates the gemini-1.5-flash-generated surprise scores, H represents human surprise scores collected through our experiments, and C shows the ground truth perceived visual complexity - all values are normalized to a 0-100 scale. The second figure illustrates cases where relatively ordinary images tend to be classified as visually complex, due to other dimensions of complexity such as high number of objects or segmentations. Figures 4 and 5 share similar visual features (number of segmentations = 88 vs. 88, number of classes = 5 vs. 4, MSG = 37 vs. 38, MUC = 34 vs. 24), yet their visual complexity values differ significantly (56 vs. 72). This difference may be attributed to how participants perceived the latter image as more surprising (76 vs. 98).
Ablations on LLMs for Surprise Scores
To generate surprise scores for images in our SVG dataset, we evaluated three different LLMs: gemini-1.5-flash (?, ?), Llama-3.2-8B (?, ?), and ChatGPT-4o (?, ?). Our selection strategy aimed to compare models across the accessibility spectrum: Llama-3.2-8B as a fully open-source model with unrestricted access, ChatGPT-4o as a sophisticated commercial model available only through paid API (or free account using the UI), and gemini-1.5-flash as a middle-ground solution offering free access with daily request limitations.
We input the exact same Chain-of-Thought prompt (as given in Algorithm 1) to the models. Although a popular choice as being fully open-source, Llama-3.2-8B could not keep up with the other models, as it struggled with understanding the task and providing responses that could be parsed by our algorithm. Therefore, we decided to focus on comparing gemini-1.5-flash and ChatGPT-4o for our task. In Table 3, we present the Spearman correlation coefficients between our predicted values and ground truth perceived visual complexity, comparing surprise scores derived from gemini-1.5-flash, ChatGPT-4o, and human participants as features in our regression models.
Table 3: Correlations of our models with the ground truth perceived visual complexity.
Surprise scores
Other features included ChatGPT-4o gemini-1.5-flash Human-rated
$\sqrt{\textit{num\_seg}}+\sqrt{\textit{num\_class}}$
0.81 0.83 0.85
$\sqrt{\textit{num\_seg}}+\sqrt{\textit{num\_class}}+\text{MSG}+\text{MUC}$
0.82
0.85
0.86
As anticipated, human-rated surprise scores yield the strongest contribution to the baseline model, enhancing correlation from 0.78 to 0.86. gemini-1.5-flash and ChatGPT-4o-generated surprisal scores show substantial agreement with a correlation of 0.78. However, the gemini-1.5-flash-generated surprise scores demonstrate superior performance compared to those produced by ChatGPT-4o, with the latter failing to surpass Feng’s supervised model (0.83). Based on these findings, we ultimately selected gemini-1.5-flash-produced surprise scores for our final model—a choice that offered the additional practical advantage of free access within certain usage limits.
Further Details on MSG and MUC
Here we present two examples in Figure 10 from the previous study (?, ?), presented as failure modes.
MSG applies the Sobel operator to each $k\times k$ patch in the image. While the Sobel operator has symmetric magnitude, its kernel contains opposite signs on different sides. Consequently, symmetrical image patches typically produce lower MSG values. This image has a high segmentation count (538 - maximum in VISC dataset), yielding a high baseline prediction of 86 versus the ground truth of 56. Incorporating its MSG score of 0.26 (50th percentile) reduces the predicted complexity to 0.69, aligning closer to the actual value.
Similarly, this image receives a high baseline prediction (57) due to its substantial segmentation count (204, 78.5 percentile) and class diversity (37, 87 percentile). However, by incorporating both symmetry effects and color uniformity, we can refine our prediction. The image’s low MSG value (0.19, 15th percentile) and MUC value (0.18 for 7 bits, 7th percentile) adjust the final prediction to 0.20, much closer to the ground truth complexity of 0.16. Both the symmetrical patches and monochromatic structure reduce its perceived visual complexity.
Figure 10: Examples where MSG and MUC features help improve complexity predictions: (top) an image from VISC dataset and (bottom) an image from Sav. Int. dataset, both previously shared by (?, ?) as failure modes.
We performed multiple ablation studies across all 16 datasets to determine the superior methodology. Our statistical analysis assessed whether the correlation between feature $X$ and visual complexity differs significantly from that between feature $Y$ and complexity, allowing us to identify the more effective approach.
The permutation test first calculates Spearman’s rank correlations ($r_{s}$) between complexity $C$ and both features $X$ and $Y$, computing the difference of their absolute values, denoted as $\Delta_{\text{obs}}=|r_{s}(C,X)|-|r_{s}(C,Y)|$. It then performs $n=1000$ permutations, each time randomly shuffling the complexity values $C^{(i)}$ while keeping feature values fixed, which breaks any real relationships between complexity and the features. For each permutation $i$, it calculates new correlations with the shuffled complexity values and records their permuted difference as $\Delta^{(i)}_{\text{perm}}=|r_{s}(C^{(i)},X)|-|r_{s}(C^{(i)},Y)|$. Statistical significance is determined by calculating a p-value:
$p=\frac{\sum_{i=1}^{n}\mathbb{I}(|\Delta^{(i)}_{\text{perm}}|\geq|\Delta_{\text{obs}}|)+1}{n+1}$
where $\mathbb{I}(\cdot)$ is the indicator function that equals 1 when the condition is true and 0 otherwise.A low $p$-value ($p<0.05$) indicates the difference between correlations is statistically significant, and if significant, the feature with the higher absolute correlation is declared the “winner”.
We carried out a preliminary small-scale ablation study on weights and scales for MSG/MUC algorithms. Our baseline used scales $[1,2,4,8]$ with weights $[0.4,0.3,0.2,0.1]$ (baseline). We compared this against three alternatives: (1) same scales with reversed weights $[0.1,0.2,0.3,0.4]$ (alternative #1); (2) fewer scales $[1,4]$ with weights $[0.6,0.4]$ (alternative #2); (3) linearly incremental scales $[1,2,3,4]$ with baseline weights (alternative #3) and (4) larger scales $[8,16,32,64]$ with baseline weights (alternative #4). For MSG, we also evaluated using a larger $5\times 5$ kernel size. For MSG, each comparison is carried out 16 times (number of datasets). For MUC, we carried out the comparison separately for each bit setting $\in\{3,4,5,6,7,8\}$ and reported the total number of win/loss/non-significant ratio ($16\times 6=96$ runs).
Feature Comparison Result (w/l/ns)
MSG baseline vs. alt. #1 7/4/5
baseline vs. alt. #2 4/3/9
baseline vs. alt. #3 4/4/8
baseline vs. alt. #4 10/1/5
baseline vs. alt. kernel 6/3/7
MUC baseline vs. alt. #1 19/17/60
baseline vs. alt. #2 22/10/64
baseline vs. alt. #3 13/8/75
baseline vs. alt. #4 51/9/36
Table 4: Results of our preliminary ablation studies. Numbers indicate datasets where the first feature significantly outperformed (win), underperformed (loss), or showed no significant difference (non-significant) compared to the second feature ($p<0.05$).
Based on our ablation study, we selected the baseline weight and scale configuration for MSG and MUC algorithms to conduct subsequent experiments. While our ablation study guided the selection of a baseline configuration, we recognize that many other parameter combinations could be explored. A more heavily tuned setup might yield even better results, but we opted for a simpler baseline to focus on the broader methodological implications without spending excessive time on hyperparameter optimization.
We then extended our investigation by examining various algorithmic variations, including colored versus grayscale MSG, single-scale versus multi-scale implementations, and comparative analyses of different feature extraction techniques. These comparisons encompassed edge density and patch-based symmetry methods previously introduced by Davidson et al. (?, ?). Table 5 highlights the effectiveness of colored MSG, multi-scale techniques, and MSG over the baseline alternatives.
Comparison Result (w/l/ns)
Colored MSG vs. Grayscale MSG 6/2/8
Multi-scale MSG vs. Single-scale MSG 8/2/6
Multi-scale MUC vs. Single-scale MUC 5/1/10
MSG vs. Edge Density 5/2/9
MSG vs. Patch-symmetry 7/1/8
Table 5: Results of feature comparison ablation studies. Numbers indicate datasets where the first feature significantly outperformed (win), underperformed (loss), or showed no significant difference (non-significant) compared to the second feature ($p<0.05$).
Conclusion
In this work, we introduce novel visual features that significantly outperform those traditionally employed in visual complexity literature, enhancing the existing segmentation-based model while preserving its interpretability. Drawing inspiration from the familiarity bias, we explore a previously unexamined dimension of visual complexity: the element of surprise. We first developed the Surprising Visual Genome (SVG) dataset, created based on the well-studied Visual Genome collection, to systematically investigate semantic factors in complexity perception. Using this resource, we conducted human participant studies that revealed a consistent tendency to perceive unfamiliar, surprising, or bizarre images as visually more complex. We then used LLMs to automate the generation of surprise scores with explicit reasoning capabilities, creating a scalable approach to quantifying this subjective dimension of visual complexity without sacrificing interpretability. Our final model—which integrates low-level visual features, segmentation and class counts, and LLM-generated surprise scores—not only outperforms all previous handcrafted feature approaches but also achieves competitive results against the best-supervised model while maintaining interpretability.